OpenStudy (haleya):

can someone help me please:)!!!!! 1. sqrt5 + 6 sqrt5

OpenStudy (solomonzelman):

If you were to add, \(x+6x\) what would you get?

OpenStudy (solomonzelman):

(By the way, I`m assuming that your question is \(\sqrt{5}+6\sqrt{5}\))

OpenStudy (haleya):

yes!um you would get 7x

OpenStudy (solomonzelman):

OK, and when you add \(\sqrt{5}+6\sqrt{5}\) (treat \(\sqrt{5}\) as if it were \(x\))

OpenStudy (haleya):

okayy

OpenStudy (haleya):

i still dont knwo how to get the answer. the answers are 6 sqrt5, 30, 6sqrt10, and 7sqrt5